OpenStudy (anonymous):

What is an equation of a parabola with the given vertex and focus? Vertex: (-2, 5) Focus: (-2,6)

OpenStudy (anonymous):

The vertex and focus are vertically aligned, so the parabola is vertical. The focus lies above the vertex, so the parabola opens upwards. General equation of up-opening parabola: y = a(x - h)² + k vertex (h, k) focal length p = 1/(4a) focus (h, k+p) Applying your data: p = 1 a = 1/(4p) = ¼ y = (¼)(x + 2)² + 5

OpenStudy (anonymous):

Thank You !

OpenStudy (anonymous):

np